# Efficient Quantum Circuit Design with a Standard Cell Approach

### 13 Jan 2024 v2

@@ Abstract
We design quantum circuits by using the standard cell approach borrowed from classical circuit design, which can speed-up the layout of circuits with a regular structure. Our standard cells are general and can be used for all types of quantum circuits: error-corrected or not. The standard cell approach enable the formulation of layout-aware routing algorithms. Herein, we use cubic standard cells for Toffoli gates and, starting from a 3D architecture, we design a multiplication circuit. We present evidence that, when compared with automatic routing methods, our layout-aware routers are significantly faster and achieve shallower 3D circuits (by at least 2.5x) and with a lower routing cost. Additionally, our co-design approach can be used to estimate the resources necessary for a quantum computation without using complex compilation methods. We conclude that standard cells, with the support of layout-aware routing, pave the way to very large scale methods for quantum circuit compilation.
@@ 1 Introduction
The qubit layout of a quantum computer is, in general, very regular and similar to a tiling. For example, NISQ devices are tilings of various polygons such as squares (Google Sycamore), octagons and squares (Rigetti Aspen), or hexagons (IBM Hummingbird). Most of the devices are 2D, but there also exist 3D proposals such as for neutral atoms [1].
This paper is organised as follows: Section 1.1 introduces the goals of our approach, and Section 1.2 connects and differentiates our method from the state-of-the-art. Section 2 is describing the standard cells, and the details about how these are used for compilation (Section 2.1). Section 2.2 shows how the tiles are helping and speeding up the routing procedures. Section 2.3 will illustrate by numbers the advantage of the tiles we are using in this paper. Finally, we compare tiling with automatic compilation and the results are presented in Section 3.
@@ 1.2 Related Work
Given the extreme complexity of VLSI circuit design, the use of standard cells have become a mainstream technique for the efficient design of large-scale high-performance computing systems [6] within standard circuit design curricula [7]. In the conventional VLSI standard cell approach, a library of optimized circuit blocks are interconnected as tiles within a large computing system [8]. This is in contrast to full-custom design, in which each transistor is individually designed in a fully custom manner; while full-custom design permits full optimization to meet design specifications, it is intractable when billions or trillions of transistors are considered. While the standard cells themselves – and the heuristics to design VLSI circuits with them – are valuable proprietary intellectual property for the large chipmakers (Intel, AMD, etc.) [9], open source options are available for academic and small-scale applications. Standard cell VLSI approaches thus enable the practical and economical design of efficient circuits with hundreds of trillions of transistors [10] with a manageable sacrifice in optimality.
@@ 1.3 Contributions
We introduce standard quantum cells, in the following also called tiles. Our method starts from the observation that qubit lattices, as well as the quantum circuit that will be compiled, are very regular structures (e.g. Fig. 2). We join tiles into designs that: a) represent the structure of the circuit to be compiled and, at the same time, b) are compatible with the computer’s qubit arrangement. Standard cell design approaches are suitable for the design of near-optimal quantum computing systems within reasonable design time constraints. Our work focuses on 3D lattices of qubits.
Our cells are designed (pre-programmed) for computations where no SWAP gates are necessary within the cell. For example, no SWAPs are required in order to perform the Toffoli gate on the cell from Fig. 1. After tiling multiple cells, qubits will still need to be routed from one tile to the other: long range interactions between remote qubits are necessary for implementing the overall computation. In this paper, we express the cost of the qubit routes through the count and depth of the nearest-neighbour SWAP circuits necessary to implement the routing.
@@ 2.3 Low-Depth Circuits by Layout-Aware Routing
In the listings, “upper” and “lower” refer to the top and bottom of the Toffoli-cube being worked in respectively; e.g. “lower W” refers to the ancillary qubit in the West corner of the bottom 4 corners of the current Toffoli-cube. “N” is for North. Multiple gates listed on a line are in the same moment (a moment is taken from the Cirq terminology, and represents all the gates that can be executed in parallel).
@@
Repeat for each a and b
    Toffoli (ctrl, $A_{0}$, $B_{0}$)
    SWAP (ctrl, W), SWAP($A_{0}$, N)
    SWAP (ctrl, $B_{0}$), SWAP ($A_{0}$, N)
    SWAP($B_{0}$, ancilla directly above)
Until just before the nth cube.
Toffoli(ctrl, $A_{n-1}$, $B_{n-1}$)
SWAP($A_{n-1}$, S), SWAP($B_{n-1}$, $B_{n}$)
SWAP(W, z), SWAP($B_{n}$, lower W), SWAP($A_{n-1}$, ctrl)
SWAP($B_{n}$, z), SWAP($B_{n-1}$, lower N)
SWAP($B_{n}$, empty space in queue), SWAP($B_{n-1}$, upper N)
SWAP($B_{n-1}$, N above current position), SWAP(z, lower N), SWAP(ctrl, W)
SWAP(ctrl, next ancilla (B qubit) in queue))
@@ 3 Results: Fast Compilation, Scalable and Efficient Routing
Fig. 5 compares the SWAP counts and SWAP depths of the circuits compiled with Cirq and our tiled versions. The compilation of the 6-qubit multiplier required more than 3 days. We repeated the compilation multiple times and measured the same duration. We did not succeed to compile using Cirq 7-qubit wide multipliers automatically within reasonable time (one week).
When compared with automatic routing methods, such as the ones available in Google Cirq, our tiled version of the multiplier improves significantly on both SWAP depth and count. Fig. 5 illustrates the comparison between the counts and depths of the manually tiled circuit and the routing costs computed automatically using Cirq. The SWAP depth and count numbers for Cirq were obtained by providing the router with a 3D device of the same general dimensions as our tiled circuit and allowing it to route greedily. In many cases the router was actually provided with more qubits than the tiled circuit uses. For additional details and source code see the repository at [29].
@@ 4 Conclusion
There exist multiple applications of standard cells and tiling. First, tiling quantum circuits can inform the co-design of computing architectures, where the qubit layout, for example, is developed in parallel to the circuits to execute. Such 2D and 3D architectural co-design can be implemented with neutral atoms [1], for example.
@@ 5 Acknowledgments
Toffoli(ctrl, $A_{3}$, z)
Toffoli($B_{j}$, $A_{0}$, $A_{1}$)
Toffoli($B_{j+1}$, $A_{1}$, $A_{2}$)
Toffoli($B_{j+2}$, $A_{2}$, $A_{3}$)
SWAP($B_{j+3}$, ctrl)
Toffoli($B_{j+3}$, $A_{3}$, z)
SWAP($B_{j+3}$, ctrl), SWAP($A_{3}$, N)
SWAP($A_{3}$, z)
SWAP(z, E), SWAP($A_{3}$, $B_{j+4}$)
Toffoli(ctrl, z, $B_{j+4}$)
SWAP($B_{j+3}$, ctrl), SWAP($A_{3}$, $B_{j+4}$), SWAP(z, N)
SWAP(z, $A_{3}$)
SWAP($A_{3}$, E)
Toffoli($B_{j+3}$, $A_{3}$, z)
SWAP(z, lower W)
SWAP($B_{j+3}$, z)
SWAP(z, ctrl), SWAP($B_{j+3}$, lower N)
for(k = n-1, k >=0; k–)
    Toffoli(ctrl, $A_{k}$, $B_{j+k}$)
    Toffoli($B_{j+k-1}$, $A_{k}-1$, $A_{k}$)
    SWAP(ctrl, next clockwise ancilla), SWAP($B_{j+k-1}$, next clockwise ancilla)
    SWAP(ctrl, $A_{k}$)
    SWAP(ctrl, $B_{j+k-1}$), SWAP($A_{k}$, next counter-clockwise ancilla)
    SWAP(ctrl, next counter-clockwise ancilla)
Toffoli(ctrl, $A_{0}$, $B_{j}$)
@@
1.  SWAP(ctrl, queue above)
    SWAP($B_{lowest}$, queue above)
    for each cube k (besides the first):
        //this is all done in the same moment
        SWAP($A_{k}$, next counter-clockwise ancilla)
@@ 1.  SWAP(ctrl, queue above)
2.  for each cube k (besides the first):
        //this is all done in the same moment
        SWAP($A_{k}$, next counter-clockwise ancilla)
        SWAP($B_{j+k}$, $ancillA_{above}$)
    SWAP(z, ancilla in the control position)
    SWAP($B_{j+n}$, N) //the ancilla above the next b in the queue
3.  SWAP($B_{j+n}$, W)
4.  SWAP($B_{j+n}$, Z)
5.  SWAP(Z, N), SWAP($B_{j+n}$, B_${j+1}$)

### 08 Apr 2024 v3

@@ Abstract
We design quantum circuits by using the standard cell approach borrowed from classical circuit design, which can speed-up the layout of circuits with a regular structure. Our standard cells are general and can be used for all types of quantum circuits: error-corrected or not. The standard cell approach enables the formulation of layout-aware routing algorithms. Our method is directly applicable to neutral atom quantum computers supporting qubit shuttling. Such computers enable zoned architectures for memory, processing and measurement, and we design circuits using qubit storages (memory and measurement zones) and standard cells (processing zones). Herein, we use cubic standard cells for Toffoli gates and, starting from a 3D architecture, we design a multiplication circuit. We present evidence that, when compared with automatic routing methods, our layout-aware routers are significantly faster and achieve shallower 3D circuits (by at least 2.5x) and with a lower routing cost. Additionally, our co-design approach can be used to estimate the resources necessary for a quantum computation without using complex compilation methods. We conclude that standard cells, with the support of layout-aware routing, pave the way to very large scale methods for quantum circuit compilation.
@@ 1 Introduction
The optimal design of large-scale circuits – both quantum and classical – is not computationally tractable, necessitating the use of sub-optimal heuristics. Noting that the qubit layout of a quantum computer is generally regular and similar to a tiling, large-scale quantum circuit design challenges can be naturally mapped to classical circuit design heuristics based on standard cells. We therefore propose the development of automated quantum circuit design algorithms that leverage the enormous repository developed through several decades of classical circuit design.
This paper is organised as follows: Section 1.1 introduces the goals of our approach, and Section 1.2 connects and differentiates our method from the state-of-the-art. Section 1.5 describes the natural application of our approach to the scalable compilation of neutral atom quantum circuits. Section 2 is describing the standard cells, and the details about how these are used for compilation (Section 2.1). Section 2.2 shows how the tiles are helping and speeding up the routing procedures. Section 2.3 will illustrate by numbers the advantage of the tiles we are using in this paper. Finally, we compare tiling with automatic compilation and the results are presented in Section 3.
@@ 1.2 Related Work
The naive “full-custom” approach for optimizing large-scale classical circuit design automation is to automatically explore the entire range of transistor parameters and all possible interconnected circuit structures that achieve a desired complex computational function. However, this is intractable when billions or trillions of transistors are considered. Therefore, while full-custom design permits full optimization to meet design specifications, it is neither computationally nor economically viable for large system designs.
Given the extreme complexity of VLSI circuit design, the use of “standard cells” have therefore become a mainstream technique for the efficient design of large-scale high-performance computing systems [5] within standard circuit design curricula [6]. In the conventional VLSI standard cell approach, a library of standard cells are developed that each contain multiple transistors with optimized parameters within an optimized circuit structure for a particular logical function. An efficient large computing system can then be automatically designed through the near-optimal interconnection of these standard cells [7]. This approach thus divides intractable large-scale design challenges into two levels of hierarchy that are both tractable, providing a computationally feasibly approach to the automated design of nearly-optimal large scale circuits.
While the standard cells themselves – and the heuristics to design VLSI circuits with them – are valuable proprietary intellectual property for the large chipmakers (Intel, AMD, etc.) [8], open source options are available for academic and small-scale applications. Standard cell VLSI approaches thus enable the practical and economical design of efficient circuits with hundreds of trillions of transistors [9] with a manageable sacrifice in optimality.
1.3 Proposed Approach
In this work, we note the fact that quantum circuit tiles are naturally analogous to standard cells, and therefore propose that large-scale quantum circuit design apply a standard cell approach adapted from classical VLSI. We particularly note that NISQ devices are tilings of various polygons such as squares (Google Sycamore), octagons and squares (Rigetti Aspen), or hexagons (IBM Hummingbird). Most of the devices are 2D, but there also exist 3D proposals such as for neutral atoms [10].
@@ 1.4 Contributions
We introduce standard quantum cells, in the following also called tiles. Our method starts from the observation that qubit lattices, as well as the quantum circuit that will be compiled, are very regular structures (e.g. Fig. 2). We join tiles into designs that: a) represent the structure of the circuit to be compiled and, at the same time, b) are compatible with the computer’s qubit arrangement.
Standard cell design approaches are suitable for the design of near-optimal quantum computing systems within reasonable design time constraints. Our work focuses on 3D lattices of qubits. Such lattices can be easily implemented by neutral atom computers, as discussed in the following section.
Our cells are designed (pre-programmed) for computations where ideally no SWAP gates are necessary within the cell. For example, no SWAPs are required in order to perform the Toffoli gate on the cell from Fig. 1. After tiling multiple cells, qubits will still need to be routed from one tile to the other: long range interactions between remote qubits are necessary for implementing the overall computation. Consequently, our approach to compilation is the repeated execution of the following two steps:
  1. 1.
    Routing of qubits between the cells;
  2. 2.
    Execution of gates within the cell.
In this paper, we express the cost of the qubit routes through the count and depth of the nearest-neighbour SWAP circuits necessary to implement the routing. Nevertheless, in the context of neutral atom computers, the SWAP cost can be replaced by the cost of shuttling the qubits.
@@
1.5 Application to Neutral Atom Quantum Computers
Neutral-atom quantum computers have demonstrated the ability to operate on thousands of qubits (e.g. [23]). These qubits are typically arranged in planar 2D layouts, while 3D configurations are feasible but more challenging to engineer. Notably, neutral-atom systems enable a technique called qubit shuttling which allows for all-to-all connectivity for two-qubit gates under certain constraints. This effectively emulates 3D architectures even within a 2D layout.
Recent experiments with neutral-atom devices (e.g. [24]) have spurred interest in scalable compilation methods that exploit qubit shuttling to reduce reliance on SWAP gates. While functionally equivalent (e.g. [25, 26]), shuttling can impact circuit execution speed and fidelity. Trade-off analyses suggest that shuttling is a viable substitute for low-fidelity SWAP gates, but not yet a replacement for high-fidelity ones [26].
One promising model is the dynamically field-programmable qubit array (DPQA) [27], which allows for reconfigurable qubit layouts during computation. The DPQA architecture utilizes mobile traps (acusto-optic deflectors, AODs) and fixed traps (spatial light modulators, SLMs) to facilitate qubit movement between them. Two-qubit gates are performed by shuttling qubits into proximity using AODs, followed by simultaneous application of laser pulses across the entire 2D plane. This multi-step approach in DPQA mirrors our proposed route-and-execute method. Computations within the DPQA model involve a three-step loop [26]: (1) loading qubits into AODs, (2) shuttling and unloading qubits into SLMs, and (3) gate execution.
Neutral-atom computers, particularly those employing AODs and SLMs, enable architectures with designated zones for specific functions [24]. These zones include memory (storing unused qubits), execution (performing gate operations), and measurement. Zone-based architectures have demonstrably facilitated pipelining in neutral-atom computations [24]. To our knowledge, our approach is the first to leverage qubit storage for co-design and compilation of circuits. This approach enables analysis and optimization of circuits as queuing networks, drawing upon prior work by Dobbs (2021).
Within the framework of specialized zones, standard cells represent the execution zones. Our approach automates the pipelining of circuit execution across sequences of zones (tiles), as detailed in Section 2.1 and Figure 4. Furthermore, our cell design allows for the pre-computation of optimal shuttling routes, leading to more scalable compilation compared to existing methods like the hybrid-SMT approach [27], hybrid heuristics [28], and qubit mapping methods [29].
@@ 2.3 Low-Depth Circuits by Layout-Aware Routing
In the listings, “upper” and “lower” refer to the top and bottom of the Toffoli-cube being worked in respectively; e.g. “lower W” refers to the ancillary qubit in the West corner of the bottom four corners of the current Toffoli-cube. “N” is for North. Multiple gates listed on a line are in the same moment (a moment is taken from the Cirq terminology, and represents all the gates that can be executed in parallel).
@@
01: Repeat from i = 0 to i = n-2
02:    Toffoli (ctrl, $A_{i}$, $B_{i}$)
03:    SWAP (ctrl, W), SWAP($A_{i}$, N)
04:    SWAP (ctrl, $B_{i}$), SWAP ($A_{i}$, N)
05:    SWAP($B_{i}$, ancilla directly above)
06: It remains to apply the Toffoli on the final cube and clean up:
07: Toffoli(ctrl, $A_{n-1}$, $B_{n-1}$)
08: SWAP($A_{n-1}$, S), SWAP($B_{n-1}$, $B_{n}$)
09: SWAP(W, z), SWAP($B_{n}$, lower W),
       SWAP($A_{n-1}$, ctrl)
10: SWAP($B_{n}$, z), SWAP($B_{n-1}$, lower N)
11: SWAP($B_{n}$, empty space in queue),
       SWAP($B_{n-1}$, upper N)
12: SWAP($B_{n-1}$, N above current position),
       SWAP(z, lower N), SWAP(ctrl, W)
13: SWAP(ctrl, next ancilla (B) in queue))
@@ 3 Results: Fast Compilation, Scalable and Efficient Routing
Fig. 5 compares the SWAP counts and SWAP depths of the circuits compiled with Cirq and our tiled versions. When compared with automatic routing methods, such as the ones available in Google Cirq, our tiled version of the multiplier improves significantly on both SWAP depth and count.
Fig. 5 illustrates the comparison between the counts and depths of the manually tiled circuit and the routing costs computed automatically using Cirq. The SWAP depth and count numbers for Cirq were obtained by providing the Cirq router with a 3D device of the same general dimensions as our tiled circuit and allowing it to route greedily. This is an extremely challenging problem, because the router needs to map the circuit qubits to the device qubits in a way that allows for non-overlapping routes in an efficient manner (lower depth). To simplify the mapping and the routing, in many cases the router was actually provided with more qubits than the tiled circuit uses. For additional details and source code see the repository at [36].
The compilation of the 6-qubit multiplier required more than 3 days. We repeated the compilation multiple times and measured the same duration. We did not succeed to compile using Cirq 7-qubit wide multipliers automatically within reasonable time (one week).
@@ 4 Conclusion
There exist multiple applications of standard cells and tiling. First, tiling quantum circuits can inform the co-design of computing architectures, where the qubit layout, for example, is developed in parallel to the circuits to execute. Such 2D and 3D architectural co-design can be implemented with neutral atoms [10], for example. In neutral atom computers, qubits can be easily shuttled and moved between specialized zones (memory, execution, measurement). Standard cells are analogous to execution zones, and by tiling cells one is designing a pipelined architecture, where qubits are moved/shuttled between zones. This allows to determine optimal shuttling schedules for the movement of the qubits between the zones.
@@ 5 Acknowledgments
01: Toffoli(ctrl, $A_{3}$, z)
02: Toffoli($B_{j}$, $A_{0}$, $A_{1}$)
03: Toffoli($B_{j+1}$, $A_{1}$, $A_{2}$)
04: Toffoli($B_{j+2}$, $A_{2}$, $A_{3}$)
05: SWAP($B_{j+3}$, ctrl)
06: Toffoli($B_{j+3}$, $A_{3}$, z)
07: SWAP($B_{j+3}$, ctrl), SWAP($A_{3}$, N)
08: SWAP($A_{3}$, z)
09: SWAP(z, E), SWAP($A_{3}$, $B_{j+4}$)
10: Toffoli(ctrl, z, $B_{j+4}$)
11: SWAP($B_{j+3}$, ctrl),
        SWAP($A_{3}$, $B_{j+4}$), SWAP(z, N)
12: SWAP(z, $A_{3}$)
13: SWAP($A_{3}$, E)
14: Toffoli($B_{j+3}$, $A_{3}$, z)
15: SWAP(z, lower W)
16: SWAP($B_{j+3}$, z)
17: SWAP(z, ctrl), SWAP($B_{j+3}$, lower N)
18: for(k = n-1, k >=0; k–)
19:     Toffoli(ctrl, $A_{k}$, $B_{j+k}$)
20:     Toffoli($B_{j+k-1}$, $A_{k}-1$, $A_{k}$)
21:     SWAP(ctrl, next clockwise ancilla),
           SWAP($B_{j+k-1}$, next clockwise ancilla)
22:     SWAP(ctrl, $A_{k}$)
23:     SWAP(ctrl, $B_{j+k-1}$),
           SWAP($A_{k}$, next counter-clockwise ancilla)
24:     SWAP(ctrl, next counter-clockwise ancilla)
25: Toffoli(ctrl, $A_{0}$, $B_{j}$)
@@
01:  SWAP(ctrl, queue above)
02:  SWAP($B_{lowest}$, queue above)
03:  for each cube k (besides the first):
04:     SWAP($A_{k}$, next counter-clockwise ancilla)
@@
05:  for each cube k (besides the first):
06:     SWAP($A_{k}$, next counter-clockwise ancilla),
            SWAP($B_{j+k}$, $ancillA_{above}$),
            SWAP(z, ancilla in control position)
07:  SWAP($B_{j+n}$, N)
08:  SWAP($B_{j+n}$, W)
09:  SWAP($B_{j+n}$, Z)
10:  SWAP(Z, N), SWAP($B_{j+n}$, B_${j+1}$)